OpenStudy (anonymous):

Which is the simplified form of 6 times x minus 18 all over 5 times x minus 15.? negative 6 over 5. x minus 3 over x minus 1. 3 over 5. 6 over 5.

OpenStudy (anonymous):

Is A correct?

OpenStudy (whpalmer4):

No.

OpenStudy (anonymous):

I remember doing this a lot but I totally forgot the formula and I'm trying to figure it out lol.

OpenStudy (whpalmer4):

Factor the numerator. Factor the denominator. What do you get?

OpenStudy (anonymous):

B

OpenStudy (whpalmer4):

Back to the old guessing game, eh? Your life in math classes is going to be miserable if you don't learn how to do this.

OpenStudy (whpalmer4):

Is there a common factor in the two terms in \[6x-18\]

OpenStudy (anonymous):

2

OpenStudy (whpalmer4):

What do you get if you divide both terms by 2?

OpenStudy (whpalmer4):

18/2 = 6?

OpenStudy (anonymous):

my bad. I meant 9

OpenStudy (whpalmer4):

So that gives us \[2(3x-9)\]Are there any common factors in \(3x\) and \(9\)?

OpenStudy (anonymous):

3?

OpenStudy (whpalmer4):

You tell me :-)

OpenStudy (anonymous):

3

OpenStudy (whpalmer4):

So what do you get if you divide \(3x\) by \(3\)? How about \(9\) by \(3\)?

OpenStudy (anonymous):

3x/3 is 1x. and 9/3=3

OpenStudy (whpalmer4):

Excellent. So that means our numerator can be written as \[2*3(1x-3) = 6(x-3)\] Now how about the denominator? \[5x-15\]See any common factors there?

OpenStudy (anonymous):

5

OpenStudy (whpalmer4):

Okay, if we factor out a 5, what do we have left?

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

No, I mean if we divide \[5x-15\]by \(5\), what do we get?

OpenStudy (anonymous):

-2x

OpenStudy (whpalmer4):

\[\frac{5x-15}5=\frac{5x}{5}-\frac{15}{5}=\]

OpenStudy (anonymous):

I don't get that

OpenStudy (whpalmer4):

Okay, what's \(5x\) divided by \(5\)?

OpenStudy (anonymous):

1x

OpenStudy (whpalmer4):

Right. and \(-15\) divided by \(5\)?

OpenStudy (anonymous):

-3

OpenStudy (whpalmer4):

Excellent. So that means we can rewrite \[5x-15\] as\[5(x-3)\]because \[5(x-3) = 5*x - 5*3 = 5x-15\] So our fraction can be written \[\frac{6(x-3)}{5(x-3)}\]Any ideas about how to simplify that sucker?

OpenStudy (anonymous):

Umm... ughh... not too sure.

OpenStudy (anonymous):

you can do this ^^ xD

OpenStudy (whpalmer4):

Well, can't you cancel out the \(x-3\) terms, which appear on both the top and the bottom? What's 3/3? What's 5/5? What's 235234/235234? Aren't they all 1?

OpenStudy (anonymous):

Yeah haha

OpenStudy (whpalmer4):

As long as \(x\ne 3\), we can just cancel those two terms out: \[\frac{6(x-3)}{5(x-3)} = \frac{6\cancel{(x-3)}}{5\cancel{(x-3)}} =\frac{6}{5}\]

OpenStudy (whpalmer4):

\(x\) isn't allowed to be \(3\) in this fraction because that would involve dividing by \(0\), and that doesn't make any sense. Doesn't matter for this problem, though.

OpenStudy (anonymous):

Awesome! Thanks!

OpenStudy (whpalmer4):

So when you have to factor something like \(6x-18\), it's worth trying right off the bat dividing both numbers by the first number to see if works. If it does, that's much faster than dividing by 2, 3, etc.

OpenStudy (anonymous):

I'll keep that in mind.

OpenStudy (whpalmer4):

You can think of this factoring business as looking at a multi-car pileup and trying to figure out what the cars looked like before they collided :-)

OpenStudy (anonymous):

Haha alright, thanks!

OpenStudy (anonymous):

oh wow